OpenStudy (anonymous):

Decide what can be used to determine the equation of a line when you have two points, but no slope. a. slope-intercept form c. point-slope form b. general form d. slope formula, then use one of the given points to find the equation

OpenStudy (anonymous):

@phi

OpenStudy (austinl):

D is what I would do myself.

OpenStudy (austinl):

\[m = \frac{y_1 - y_2}{x_1 - x_2}\] \[y = mx + b\] You have (x1 , y1) , (x2 , y2) You could then plug in one of those points and solve for b, the y intercept.